Let N1=40, X1=30, N2=40 and X2=20

Calculate the test statistic ZSTAT based on the difference P1-P2

the test statistic, ZSTAT is what?
Answer #1

Solution :

Given that,

1 = x1 / n1 = 30 / 40 = 0.75

2 = x2 / n2 = 20 / 40 = 0.5

= (x1 + x2) / (n1 + n2) = (30 + 20) / (40 + 40) = 0.625

1 - = 0.375

Z = (1 - 1) / * (1 - ) (1 / n1 + 1 / n2)

Z = (0.75 - 0.50) / 0.625  * 0.375 (1 / 40 + 1 / 40)

Z = 2.309

Test statistic = 2.309